OpenStudy (r_josh10):

I really need help with some Math problems. Will fan and medal. The questions will be below.

OpenStudy (r_josh10):

Here is the question.

OpenStudy (chosenmatt):

what do you think?

OpenStudy (anonymous):

Um, we can cross multiply: 3 * (x + 2y) = 3x + 6y 5 * x = 5x So we have: 3x + 6y = 5x Subtract 6y to both sides: 3x = 5x - 6y Subtract 5x to both sides: -2x = -6y Divide 'y' to both sides: -2x/y = -6 Divide -2 to both sides: x/y = -6/-2 Divide that..not sure if it's correct..

OpenStudy (anonymous):

Finding the ratio from 'x to y' is the same thing as solving for 'x/y'

OpenStudy (r_josh10):

@iGreen. I will tell you if its correct. thankyou :)

OpenStudy (anonymous):

No problem..just divide -6 / -2..

OpenStudy (r_josh10):

The answer is 3.